OpenStudy (anonymous):

Convert the function into standard form. Show your work y=(x-4)(x+2)

OpenStudy (anonymous):

Convert the function into intercept form. Show your work y=x^2-3x-24

OpenStudy (***[isuru]***):

u want answer for 1st one or 2nd one ?

OpenStudy (anonymous):

both

OpenStudy (***[isuru]***):

for the first one... ax^2 + bx + c is the slandered form . to get that expand the expression (x-4)(x+2) wt will u get ?

OpenStudy (***[isuru]***):

y = (x-4)(x+2) = x^2 + 2x -4x -8 = x^2 -2x -8 this is the slandered form... did u get it?

OpenStudy (anonymous):

yes

OpenStudy (***[isuru]***):

now for the second one... pls check ur question again... r u sure it's x^2-3x-24 ?

OpenStudy (anonymous):

y=x^2-3x-24

OpenStudy (***[isuru]***):

to convert it into intercept forms mean u have to write it like (x + or - "something")(x + or - "something") i.e for example.. (x +3)(x-2) (x-4)(x-1) these r the examples for intercept form...

OpenStudy (***[isuru]***):

which mean u have to write y=x^2-3x-24 in this form (x + _)(x+ _) to do that we have to factorize it.. but y=x^2-3x-24 doesn't factorize easily that's the problem here ......

OpenStudy (***[isuru]***):

wt ?? wt is ur answer ?

OpenStudy (***[isuru]***):

ummm... ur reply is hard to understand lol wht did u mean by this ? "what is is converted into intercept form"

OpenStudy (anonymous):

can u answer the question? Convert the function into intercept form. Show your work y=x^2-3x-24

OpenStudy (***[isuru]***):

y=x^2-3x-24 = (x - 3/2)^2 -24 - 9/4 = (x - 3/2)^2 -105/4 this is the vertex form.. use this and find the x coordinates when y is 0 which means solve this when (x - 3/2)^2 -105/4 = 0 ( x - 3/2)^2 = 105/4 \[(x -\frac{ 3 }{ 2 }) = \frac{\pm \sqrt{105}}{ 2 }\] \[x = \frac{ 3\pm \sqrt{105} }{ 2}\] so the graph will cut the x axis when \[x = \frac{ 3 + \sqrt{105} }{2 } \ or \ x = \frac{ 3 - \sqrt{105} }{2 }\] so... the intercept form is.. \[y = ( x -\frac{ 3 +\sqrt{105} }{ 2 } )(x - \frac{ 3 - \sqrt{105} }{ 2 })\] usually we don't write this kind of intercept formm...

OpenStudy (***[isuru]***):

and ur graph is something like this.. http://fooplot.com/plot/8ge9iug37x u can c that "x" axis is cut when x is not a whole number..

OpenStudy (***[isuru]***):

that's y I asked u to double check the problem... it would have been much easier if it was y = y=x^2-2x-24